OpenStudy (anonymous):

A ring slides along a frictionless rod, and the rod itself is rotating in the horizontal plane about one end of the rod. The position of the ring on the rod measured from the fixed end of the rotating rod is R = 100 t 2 meters, while angle that the rod has rotated in time t is  = t 3 radians, where the time t is measured in seconds. Find the magnitudes the velocity and acceleration vectors of the ring relative to ground.

OpenStudy (vincent-lyon.fr):

Use: \(x(t)=r(t) \cos \theta (t)\) \(y(t)=r(t) \sin \theta (t)\)